# CVQA: Culturally-diverse Multilingual Visual Question Answering Benchmark

### v1

@@ Abstract
Visual Question Answering (VQA) is an important task in multimodal AI, and it is often used to test the ability of vision-language models to understand and reason on knowledge present in both visual and textual data. However, most of the current VQA models use datasets that are primarily focused on English and a few major world languages, with images that are typically Western-centric. While recent efforts have tried to increase the number of languages covered on VQA datasets, they still lack diversity in low-resource languages. More importantly, although these datasets often extend their linguistic range via translation or some other approaches, they usually keep images the same, resulting in narrow cultural representation. To address these limitations, we construct CVQA [^1] https://huggingface.co/datasets/afaji/cvqa, a new Culturally-diverse multilingual Visual Question Answering benchmark, designed to cover a rich set of languages and cultures, where we engage native speakers and cultural experts in the data collection process. As a result, CVQA includes culturally-driven images and questions from across 28 countries on four continents, covering 26 languages with 11 scripts, providing a total of 9k questions. We then benchmark several Multimodal Large Language Models (MLLMs) on CVQA, and show that the dataset is challenging for the current state-of-the-art models. This benchmark can serve as a probing evaluation suite for assessing the cultural capability and bias of multimodal models and hopefully encourage more research efforts toward increasing cultural awareness and linguistic diversity in this field.
Figure 1: We propose CVQA, a large-scale multilingual VQA benchmark, representing the cultures of 28 countries and 26 different languages across 10 diverse categories, comprising 9k samples.
@@ 1 Introduction
To address these limitations, we propose CVQA: a novel, large-scale, multilingual, culturally nuanced VQA benchmark that includes a diverse set of languages, including many that are underrepresented and understudied. CVQA follows the grassroots crowd-sourcing collaboration approaches taken by Masakhane for Africa [37], NusaCrowd for Indonesia [4], and AI4Bharat for India [20]. In our case, however, we collaborate across communities, rather than within one particular community, in order to maximize cultural and linguistic representation. Consequently, our data consists of 9k questions across 28 countries, covering 26 languages. We also sub-categorize CVQA based on Country-Language pairs, resulting in 33 distinct pairs, which is substantially more extensive than existing VQA benchmarks. Furthermore, each sample in CVQA falls into one of 10 diverse categories (see Table 1) and is annotated and validated by fluent speakers and those familiar with the respective cultures, ensuring high quality and diversity. Lastly, CVQA is written in both English and local languages, enabling us to benchmark multilingual MLLMs and English-only MLLMs.
@@
  • First, we introduce CVQA, a new culturally diverse multilingual visual question answering dataset consisting of over 9,000 questions from across 28 countries and 26 languages.
@@ 2.3 Data Statistics
To ensure sufficient question variation, we set the minimum number of questions to be included in CVQA to be at least 200 questions per Country-Language subset. In the end, we gathered 9,044 total questions across all subsets. Some statistics of our collected data are shown in Table 2. Our CVQA covers a diverse set of languages and locations spread across the globe. We also capture languages written in various scripts. While Latin is the dominant script (used in 20 Country-Language pairs), the remaining scripts are diverse; covering Arabic, Amharic, Bengali, Chinese, Cyrillic, Hangul, Japanese, Perso-Arabic, Sinhalese, and Tamil. The Country-Language pairs and corresponding scripts are shown in Appendix E. CVQA covers several less commonly studied languages and regions, such as Ireland-Irish, Indonesia-Minangkabau, Indonesia-Javanese, France-Breton, Nigeria-Igbo and Mongolia-Mongolian.
@@
No. of images 4,560
No. of questions 9,044
No. of countries 28
No. of languages 26
No. of Country-Language pairs 33
Avg. question per image 1.98
Avg. words per question 10.2
Avg. words per option 1.90
@@ 4 Results
The experimental results also highlight a substantial performance gap between open and closed-source MLLMs. Closed models like GPT-4o and Gemini-1.5-Flash demonstrate superior performance, with GPT-4o achieving the highest accuracy in both English (79.1%) and local language (79.4%) prompts. In contrast, open models like InstructBLIP and mBLIP-mT0 exhibit lower performance, particularly in local language prompts, indicating a need for more diverse training data and refined fine-tuning processes. While proprietary models show superior performance, it is hard to fully explain why, due to their closed nature. Additionally, their results are not reproducible. Therefore, we use open models in the rest of our experiments.
@@
We show the breakdown performances of models per category in Table 5, where the categories themselves are described in Section 2.1. Note that the category People and Everyday Life consistently achieves the highest accuracy across most models, with LLaVA-1.5-7B obtaining 58.6% in English prompts. This can be possibly attributed to the extensive training data available for everyday human activity and interaction, which widely existed in many visual-related datasets. Conversely, the Cooking & Food and Pop Culture categories exhibit lower accuracy across models, especially in local language prompts. This demonstrates that the high diversity in food and pop culture across different cultures poses a great challenge for the generalization of MLLMs.
@@
Brands 47.7 36.5 37.8 35.4 40.1 31.3 33.5 31.3 40.3 36.5 46.9 32.2
Food 45.9 32.5 33.1 29.1 40.8 30.9 27.8 26.9 37.4 30.7 43.8 32.0
Geography 46.1 37.9 36.5 35.6 41.0 32.7 31.1 32.6 35.1 32.6 43.6 32.5
Objects 50.5 34.7 36.4 32.7 39.0 24.2 32.4 30.9 42.3 31.8 52.0 31.6
People 58.6 38.7 45.7 38.2 44.9 36.2 34.4 34.1 47.1 37.6 58.3 38.1
Plants & Animals 54.8 39.4 40.6 31.9 45.9 27.9 35.4 35.9 46.4 38.2 56.8 34.8
Pop Culture 43.5 36.7 32.6 31.0 44.5 37.1 30.6 31.2 35.8 30.7 45.0 35.5
Sports 52.0 40.0 40.2 32.6 45.7 34.6 33.3 30.7 41.8 34.4 49.7 36.5
Tradition 48.8 37.1 36.6 34.9 42.4 31.8 31.6 32.1 39.0 32.9 46.5 30.5
Vehicles 50.9 43.8 38.4 39.4 40.5 32.6 37.4 35.7 42.1 33.3 52.6 33.3
@@
Self-made Image 47.0 34.0 37.9 34.0 41.5 30.4 34.4 31.9 40.5 34.2 49.1 31.1
Web Image 48.0 37.0 36.8 33.6 44.1 33.1 34.0 32.8 39.6 33.5 46.7 33.1
@@
Location-aware 48.9 36.5 36.9 33.4 42.3 31.5 31.5 31.2 39.4 33.0 47.8 32.7
Location-agnostic 47.7 35.4 37.5 33.8 43.1 31.5 34.2 32.2 39.8 33.6 47.5 32.1
The performance of various models on self-made versus web images is shown in Table 7. One of the interesting findings is the performance variability across image sources for different models. For self-made images, the accuracy of some models such as LLaVA-1.5-7B and CLIP tends to be lower compared to web images. For instance, LLaVA-1.5-7B achieves a 47.0% accuracy in English prompts on self-made images but slightly higher at 48.0% on web images. CLIP shows an accuracy of 44.1% in English prompts on web images compared to 41.5% on self-made images. While this trend is not consistent across the other models, the results still indicate that web images might be more representative of the data these models (such as LLaVA-1.5-7B and CLIP) were trained on, leading to better performance.
@@
Most of the evaluations we conduct on CVQA are under a multiple-choice setting. However, the multiple-choice setting is often brittle towards option ordering [38, 54], and not very natural with respect to real-world scenarios [30]. In this paragraph, we explore the model’s performance on CVQA in an open-ended QA setting. To evaluate in this setting, we prompt the models without giving them the options (e.g., “In which city is this monument located?”). Then, the answer is selected by choosing the model’s highest probability of generating the full answer phrase of one of the options [11] (e.g., Jakarta, Bandung, Bali, Surabaya). This way, it is robust towards ordering unlike predicting the answer letter (e.g., A), while also not giving the model multiple-choice options that can be indirectly used for deductive reasoning. Our result shows that LLaVA-1.5-7B achieved a noticeable performance drop when prompted without multiple choice, from 48.9% to just 30% average performance. This notes that in a more practical scenario, these models might be even more unreliable in cultural understanding.

### v2

@@ Abstract
Visual Question Answering (VQA) is an important task in multimodal AI, and it is often used to test the ability of vision-language models to understand and reason on knowledge present in both visual and textual data. However, most of the current VQA models use datasets that are primarily focused on English and a few major world languages, with images that are typically Western-centric. While recent efforts have tried to increase the number of languages covered on VQA datasets, they still lack diversity in low-resource languages. More importantly, although these datasets often extend their linguistic range via translation or some other approaches, they usually keep images the same, resulting in narrow cultural representation. To address these limitations, we construct CVQA [^1] https://huggingface.co/datasets/afaji/cvqa, a new Culturally-diverse multilingual Visual Question Answering benchmark, designed to cover a rich set of languages and cultures, where we engage native speakers and cultural experts in the data collection process. As a result, CVQA includes culturally-driven images and questions from across 30 countries on four continents, covering 31 languages with 13 scripts, providing a total of 10k questions. We then benchmark several Multimodal Large Language Models (MLLMs) on CVQA, and show that the dataset is challenging for the current state-of-the-art models. This benchmark can serve as a probing evaluation suite for assessing the cultural capability and bias of multimodal models and hopefully encourage more research efforts toward increasing cultural awareness and linguistic diversity in this field.
Figure 1: We propose CVQA, a large-scale multilingual VQA benchmark, representing the cultures of 30 countries and 31 different languages across 10 diverse categories, comprising 10k samples.
@@ 1 Introduction
To address these limitations, we propose CVQA: a novel, large-scale, multilingual, culturally nuanced VQA benchmark that includes a diverse set of languages, including many that are underrepresented and understudied. CVQA follows the grassroots crowd-sourcing collaboration approaches taken by Masakhane for Africa [37], NusaCrowd for Indonesia [4], and AI4Bharat for India [20]. In our case, however, we collaborate across communities, rather than within one particular community, in order to maximize cultural and linguistic representation. Consequently, our data consists of 10k questions across 30 countries, covering 31 languages. We also sub-categorize CVQA based on Country-Language pairs, resulting in 39 distinct pairs, which is substantially more extensive than existing VQA benchmarks. Furthermore, each sample in CVQA falls into one of 10 diverse categories (see Table 1) and is annotated and validated by fluent speakers and those familiar with the respective cultures, ensuring high quality and diversity. Lastly, CVQA is written in both English and local languages, enabling us to benchmark multilingual MLLMs and English-only MLLMs.
@@
  • First, we introduce CVQA, a new culturally diverse multilingual visual question answering dataset consisting of over 10,000 questions from across 30 countries and 31 languages.
@@ 2.3 Data Statistics
To ensure sufficient question variation, we set the minimum number of questions to be included in CVQA to be at least 200 questions per Country-Language subset. In the end, we gathered 10,374 total questions across all subsets. Some statistics of our collected data are shown in Table 2. Our CVQA covers a diverse set of languages and locations spread across the globe. We also capture languages written in various scripts. While Latin is the dominant script (used in 22 Country-Language pairs), the remaining scripts are diverse; covering Amharic, Arabic, Bengali, Chinese, Cyrillic, Devanagari, Hangul, Japanese, Perso-Arabic, Sinhalese, Tamil, and Telugu. The Country-Language pairs and corresponding scripts are shown in Appendix E. CVQA covers several less commonly studied languages and regions, such as Ireland-Irish, Indonesia-Minangkabau, Indonesia-Javanese, France-Breton, Nigeria-Igbo and Mongolia-Mongolian.
@@
No. of images 5,239
No. of questions 10,374
No. of countries 30
No. of languages 31
No. of country-language pairs 39
Avg. questions per image 1.98
Avg. words per question 7.6
Avg. words per option 1.80
@@ 4 Results
The experimental results also highlight a substantial performance gap between open and closed-source MLLMs. Closed models like GPT-4o and Gemini-1.5-Flash demonstrate superior performance, with GPT-4o achieving the highest accuracy in both English (75.4%) and local language (74.3%) prompts. In contrast, open models like InstructBLIP and mBLIP-mT0 exhibit lower performance, particularly in local language prompts, indicating a need for more diverse training data and refined fine-tuning processes. While proprietary models show superior performance, it is hard to fully explain why, due to their closed nature. Additionally, their results are not reproducible. Therefore, we use open models in the rest of our experiments.
@@
We show the breakdown performances of models per category in Table 5, where the categories themselves are described in Section 2.1. Note that the category People and Everyday Life consistently achieves the highest accuracy across most models, with InstructBLIP obtaining 59.8% in English prompts. This can be possibly attributed to the extensive training data available for everyday human activity and interaction, which widely existed in many visual-related datasets. Conversely, the Cooking & Food and Pop Culture categories exhibit lower accuracy across models, especially in local language prompts. This demonstrates that the high diversity in food and pop culture across different cultures poses a great challenge for the generalization of MLLMs.
@@
Brands 49.9 36.5 37.2 35.7 36.6 29.7 33.7 30.8 40.5 35.1 48.4 32.6
Food 45.4 31.9 34.5 29.1 39.2 30.4 28.1 27.6 37.7 29.8 44.4 30.6
Geography 47.1 38.2 37.1 34.2 41.8 31.9 30.6 31.6 35.0 32.3 45.3 33.2
Objects 51.8 33.0 39.4 34.5 39.7 25.4 34.3 33.0 43.1 34.0 52.3 29.1
People 58.9 38.1 45.0 37.8 46.8 30.9 35.3 34.7 46.3 36.7 59.8 34.0
Plants & Animals 55.7 37.5 43.7 32.0 48.0 27.2 35.2 35.5 46.0 36.0 55.4 35.1
Pop Culture 44.5 36.3 33.7 31.5 46.1 36.3 28.8 29.9 35.7 30.7 45.1 34.6
Sports 50.7 39.1 39.3 33.3 43.5 32.4 32.6 31.4 40.1 34.9 50.5 34.7
Tradition 50.4 35.8 37.0 35.2 41.9 32.2 31.6 31.5 39.0 32.2 47.9 30.8
Vehicles 50.6 41.4 39.5 41.1 44.6 30.5 35.6 33.9 42.0 34.0 55.0 33.0
@@
Self-made Image 48.8 34.2 38.1 34.3 41.2 30.1 31.2 31.5 40.1 33.4 48.3 31.5
Web Image 49.7 37.4 37.4 33.3 43.1 31.8 31.9 31.2 38.7 32.3 49.1 33.1
@@
Location-aware 49.6 35.5 38.0 33.7 42.7 30.6 31.3 30.9 39.3 32.7 49.0 31.9
Location-agnostic 48.3 34.7 38.1 33.8 43.8 30.8 34.1 31.8 39.8 33.6 48.7 31.1
The performance of various models on self-made versus web images is shown in Table 7. One of the interesting findings is the performance variability across image sources for different models. For self-made images, the accuracy of some models such as LLaVA-1.5-7B and CLIP tends to be lower compared to web images. For instance, LLaVA-1.5-7B achieves a 48.8% accuracy in English prompts on self-made images but slightly higher at 49.7% on web images. CLIP shows an accuracy of 43.1% in English prompts on web images compared to 41.2% on self-made images. While this trend is not consistent across the other models, the results still indicate that web images might be more representative of the data these models (such as LLaVA-1.5-7B and CLIP) were trained on, leading to better performance.
@@
Most of the evaluations we conduct on CVQA are under a multiple-choice setting. However, the multiple-choice setting is often brittle towards option ordering [38, 54], and not very natural with respect to real-world scenarios [30]. In this paragraph, we explore the model’s performance on CVQA in an open-ended QA setting. To evaluate in this setting, we prompt the models without giving them the options (e.g., “In which city is this monument located?”). Then, the answer is selected by choosing the model’s highest probability of generating the full answer phrase of one of the options [11] (e.g., Jakarta, Bandung, Bali, Surabaya). This way, it is robust towards ordering unlike predicting the answer letter (e.g., A), while also not giving the model multiple-choice options that can be indirectly used for deductive reasoning. Our result shows that LLaVA-1.5-7B achieved a noticeable performance drop when prompted without multiple choice, from 49.6% to just 30% average performance. This notes that in a more practical scenario, these models might be even more unreliable in cultural understanding.
@@ 7 Conclusion
Acknowledgments
We thank Ananjay Goel, Aditya Rachman Putra, Radityo Eko Prasojo, Amr Keleg, and Mostafa Awad for the contributions in creating the dataset. Tiago Torrent was partially funded by CNPq Research Productivity Grant number 315749/2021-0. Luis Fernando D’Haro, Mario Rodriguez-Cantelar and Marcos Estecha-Garitagoitia were supported by the European Commission through Project ASTOUND (101071191 — HORIZON-EIC-2021- PATHFINDERCHALLENGES-01), and by project BEWORD (PID2021-126061OB-C43) funded by MCIN/AEI/10.13039/501100011033 and, as appropriate, by “ERDF A way of making Europe”, by the “European Union”. We also thank the anonymous reviewers for their valuable feedback and suggestions that helped improve this paper.
@@ Appendix E Country-Language Pairs and Scripts
India Hindi Devanagari
India Marathi Devanagari
@@
India Telugu Telugu
India Urdu Perso-Arabic
@@
Appendix F Affiliation Lists
Table LABEL:tab:authors lists the authors and their respective affiliations.
Table 9: Author affiliations
Author Affiliation Author Affiliation Author Affiliation
David Romero MBZUAI Chenyang Lyu MBZUAI Haryo Akbarianto Wibowo MBZUAI
Teresa Lynn MBZUAI Injy Hamed MBZUAI Aditya Nanda Kishore IIT Madras
Aishik Mandal TU Darmstadt Alina Dragonetti Universidad de la República Artem Abzaliev University of Michigan
Atnafu Lambebo Tonja Independent Researcher Bontu Fufa Balcha Independent Researcher Chenxi Whitehouse University of Cambridge
Christian Salamea Universidad Politécnica Salesiana Dan John Velasco Samsung Research Philippines David Ifeoluwa Adelani Independent Researcher
David Le Meur Bretagne numérique Emilio Villa-Cueva MBZUAI Fajri Koto MBZUAI
Fauzan Farooqui Independent Researcher Frederico Belcavello Federal University of Juiz de Fora Ganzorig Batnasan United Arab Emirates University / MBZUAI
Gisela Vallejo The University of Melbourne Grainne Caulfield Dublin City University Guido Ivetta Universidad Nacional de Córdoba
Haiyue Song NICT Henok Biadglign Ademtew EAII Hernán Maina Universidad Nacional de Córdoba/CONICET
Holy Lovenia AI Singapore Israel Abebe Azime Saarland University Jan Christian Blaise Cruz Samsung Research Philippines
Jay Gala MBZUAI Jesus-German Ortiz-Barajas MBZUAI Jiahui Geng MBZUAI
Jinheon Baek KAIST Jocelyn Dunstan Escudero Pontificia Universidad Católica de Chile Kumaranage Ravindu Yasas Nagasinghe MBZUAI
Laura Alonso Alemany Universidad Nacional de Córdoba Luciana Benotti Universidad Nacional de Córdoba/CONICET Luis Fernando D’Haro Universidad Politecnica de Madrid
Marcelo Viridiano Federal University of Juiz de Fora Marcos Estecha-Garitagoitia Universidad Politécnica de Madrid Maria Camila Buitrago Cabrera University of Stuttgart
Mario Rodríguez-Cantelar Universidad Politécnica de Madrid Mélanie Jouitteau IKER, CNRS Mihail Mihaylov MBZUAI
Mohamed Fazli Mohamed Imam MBZUAI Muhammad Farid Adilazuarda MBZUAI Munkh-Erdene Otgonbold United Arab Emirates University
Munkhjargal Gochoo United Arab Emirates University Naome A. Etori Independent Researcher Olivier NIYOMUGISHA Independent Researcher
Paula Mónica Silva Millenium Institute Foundational Reseach on Data Pranjal Chitale Independent Researcher Raj Dabre IIT Madras
Rendi Chevi MBZUAI Ruochen Zhang Brown University Ryandito Diandaru ITB
Samuel Cahyawijaya HKUST Santiago Góngora Universidad de la República Soyeong Jeong KAIST
Sukannya Purkayastha TU Darmstadt Tatsuki Kuribayashi MBZUAI Thanmay Jayakumar IIT Madras
Tiago Timponi Torrent Federal University of Juiz de Fora, CNPq Toqeer Ehsan MBZUAI Vladimir Araujo KU Leuven
Yova Kementchedjhieva MBZUAI Zara Burzo Skyline Highschool Zheng Wei Lim The University of Melbourne
Zheng-Xin Yong Brown University Oana Ignat University of Michigan Joan Nwatu University of Michigan
Rada Mihalcea University of Michigan Thamar Solorio MBZUAI Alham Fikri Aji MBZUAI